OpenStudy (anonymous):

a right circular cylinder has a lateral area of 48 pi a radius of 4mm. find the height of the cylinder.

OpenStudy (anonymous):

If someone could like walk me through this, that'd be great!

OpenStudy (youngsuperman):

hold on may be i can help u

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

since your doing lateral area its\[ 2*r*\pi*h\] (i think)then you should be able to do it from here

OpenStudy (anonymous):

so its basically \[48 \pi=2*4 \pi*h\]

OpenStudy (anonymous):

then\[48=8*h\]

OpenStudy (anonymous):

then divide both sides by 8 to get the height of the cylinder